Sulfuryl chloride is an inorganic compound with the formula SO2Cl2. It is a colourless liquid with a pungent odour. A quantity of 5.75 g of SO2Cl2 was placed in a 3.00 L flask. At 375 °C, there is 0.035 mole of Cl2 present. SO2Cl2(g) SO2(g) + Cl2(g) .calculate kc and kp


1
Expert's answer
2021-07-20T10:50:01-0400

Molar Mass of SO2Cl2 = 134.965

5.75/134.965 = 0.0426 mole

0.0426/3 = 0.0142mole


Mole of Cl2 = 0.035

0.035/3 = 0.0117

0.0142-0.0117= 0.0025

Kc= 0.0117 × 0.0117/0.0025

= 0.0548